OpenStudy (anonymous):

How do I simplify √x^2+8x+16

OpenStudy (anonymous):

\[\sqrt{x^2+8x+16}\]

OpenStudy (unklerhaukus):

can you factorise \(x^2+8x+16\) into a perfect square \((x+a)^2\) ?

OpenStudy (unklerhaukus):

\[(x+a)^2=(x+a)(x+a)\\=x(x+a)+a(x+a)\\=x^2+ax+ax+a^2\\=x^2+2ax+a^2\]

OpenStudy (anonymous):

tells me in the book that its |x+4|

OpenStudy (unklerhaukus):

so a=4 ?

OpenStudy (anonymous):

Don't know, doesn't say what x or a are.

OpenStudy (unklerhaukus):

\[ \sqrt{x^2+2ax+a^2}=\sqrt{(x+a)^2}=|x+a|\]